A cat chases a mouse across a 1.7 m high table. The mouse steps out of the way, and the cat slides off the table and strikes the floor 1.5 m from the edge of the table.
The acceleration of gravity is 9.81 m/s2 .
What was the cat’s speed when it slid off the table?
1
Expert's answer
2019-09-25T08:37:21-0400

"\begin{cases}
L = vt; \\
H = \frac{1}{2}gt^2;
\end{cases} \\
t = \sqrt{\frac{2H}{g}}; \\
v = \frac{L}{t} = L\sqrt{\frac{g}{2H}} = 1.5\text{ m}\sqrt{\frac{9.81\frac{\text{m}}{\text{s}^2}}{2\cdot1.7\text{ m}}} \approx 2.55 \frac{\text{m}}{\text{s}}."